OpenStudy (anonymous):

give your answers as fractions -6x-3y=10 x+2y=4

OpenStudy (anonymous):

(x-3)^3

OpenStudy (anonymous):

-6x-3y=10........equation 1 x+2y=4 or x=-2y+4 we subtitute equation 2 to equation 1. therefore -6(-2y+4)-3y=10 y=34/9 x=-2(34/9)+4 x=-32/9